# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ Abstract
Directed graphs naturally model systems with asymmetric, ordered relationships, essential to applications in biology, transportation, social networks, and visual understanding. Generating such graphs enables tasks such as simulation, data augmentation and novel instance discovery; however, directed graph generation remains underexplored. We identify two key factors limiting progress in this direction: first, modeling edge directionality introduces a substantially larger dependency space, making the underlying distribution harder to learn; second, the absence of standardized benchmarks hinders rigorous evaluation. Addressing the former requires more expressive models that are sensitive to directional topologies. We propose Directo, the first generative model for directed graphs built upon the discrete flow matching framework. Our approach combines: (i) principled positional encodings tailored to asymmetric pairwise relations, (ii) a dual-attention mechanism capturing both incoming and outgoing dependencies, and (iii) a robust, discrete generative framework. To support evaluation, we introduce a benchmark suite covering synthetic and real-world datasets. It shows that our method performs strongly across diverse settings and even competes with specialized models for particular classes, such as directed acyclic graphs. Our results highlight the effectiveness and generality of our approach, establishing a solid foundation for future research in directed graph generation.
@@ 1 Introduction
Directed graphs (digraphs) naturally model systems with asymmetric relationships, capturing essential structures such as flows, dependencies, and hierarchies that arise in many real-world applications. This makes digraphs particularly well-suited for problems in diverse domains including biology li2006directed ; takane2023directed ; wei2024inference , transportation concas2022chained , social dynamics schweimer2022generating , and, more recently, image and video understanding Chang23Scenegraphs ; rodin2023actionscenegraphslongform , where structured, directional representations are critical for interpretation and reasoning. Consequently, generating digraphs is central to tasks such as such as simulation, data augmentation and novel instance discovery in domains characterized by directional structure.
@@
Figure 1: (a) The learnable space increases drastically with the number of nodes for digraphs compared to undirected graphs harary1973enumeration , highlighting challenges in extending graph generative models to directed structures. (b) Overview of our generative model for directed graphs. During training, the dual attention transformer (denoising network) parameterized by $\theta$, and enhanced with asymmetric positional encodings, learns to reverse predictions using cross-entropy loss. During inference, we compute the rate matrix $R_{t}^{\theta}(G_{t},G_{t+\Delta t})$ based on the model prediction $p^{\theta}(\cdot\mid G_{t})$, which governs the evolution of the generative process over finite intervals $\Delta t$.
@@
Graph generative models offer a data-driven approach to producing plausible and diverse samples, with applications in drug discovery mercado2020graph ; vignac2023midi , finance li2023moneylaundering , social network modeling tsai2023cdgraphdualconditionalsocial , and medicine nikolentzos2023synthetic . Despite their broad applicability and strong performance, the vast majority of existing generative models focus on undirected graphs liao2019gran ; martinkus2022SPECTRE ; vignac2023digress ; siraudin2024cometh ; xu2024disco ; Eijkelboom2024VariationalFM ; qin2025defog . Digraphs have received comparatively less attention, with recent work proposing auto-regressive models for DAGs zhang2019dvae ; li2024layerdag and, more recently, for general digraphs law2024directedheatkernels . We identify two main factors behind this limited progress. First, at the modeling level, the directed setting is inherently more challenging than its undirected counterpart, as edge directionality greatly enlarges the learnable space (see Fig.1(a): for graphs with $4$ nodes, there are $218$ possible digraphs, compared to only $11$ undirected ones). This is further amplified in settings where domain-specific structural constraints must be respected, such as directed acyclic graphs (DAGs). Second, beyond these technical challenges, we identify the lack of standardized benchmarks for directed graph generation as an important barrier to rigorous evaluation and fair comparison across methods. Current methodsli2024layerdag typically rely on proxy metrics derived from downstream task performance, without directly evaluating generated graph quality.
@@
To address the first challenge, we introduce Directo  [^1] Code available at: https://github.com/acarballocastro/DIRECTO, the first iterative refinement-based generative model for directed graphs. Our approach brings together three core components of expressive graph generative modeling: (i) informative input features, (ii) an expressive neural architecture, and (iii) a robust generative framework. For input features, we find that, upon comprehensive evaluation, positional encodings specifically designed for graphs with an asymmetric adjacency matrix, are more successful than directionality-agnostic baselines. For the architecture, we employ a dual attention mechanism that performs cross-attention between edge features and their reversed counterparts, capturing both source-to-target and target-to-source information flow. Finally, these components are integrated within a robust discrete-state flow matching framework that enables efficient generation with state-of-the-art performance. Figure 1(b) gives an overview of our method and its components.
@@
To tackle the second challenge, we propose a benchmarking framework for rigorous evaluation of directed graph generation. It spans both synthetic and real-world datasets and systematic evaluation metrics tailored to directed graphs. The synthetic benchmarks cover different distributions, including acyclic and community-based graphs, while real-world benchmarks include a DAG-focused dataset for Neural Architecture Search (NAS) google2023tpugraphs and a scene understanding dataset krishna2017visualgenome . We evaluate Directo on the proposed benchmarks, where it consistently achieves strong performance across all settings, underscoring its effectiveness across structurally diverse datasets. Extensive ablations further reveal that the proposed dual attention mechanism is critical for modeling directional dependencies, while positional encodings enhance overall generation quality.
@@
    We propose Directo, the first flow-based generative model specifically tailored for directed graphs, combining a direction-aware dual attention block with positional encoding tailored to capture directionally asymmetric structural properties.
@@
    We address the lack of benchmarks for directed graph generation by releasing a unified evaluation pipeline that encompasses both synthetic graphs with important structural properties and real-world datasets with meaningful applications.
@@
    We conduct an extensive empirical analysis demonstrating that Directo achieves state-of-the-art performance on structurally diverse synthetic and real-world graphs.
Overall, by addressing the two main bottlenecks in directed graph generation, we aim to lay the groundwork for future research and practical deployment of directed graph generation.
@@ 2 Background: Discrete Flow Matching for Graph Generation
We denote by $G=\left(x^{(1:n:N)},\,e^{(1:i\neq j:N)}\right)$ directed graphs with $N$ nodes. The set of nodes is denoted by $\{x^{(n)}\}_{n=1}^{N}$, and the set of directed edges by $\{e^{(i,j)}\}_{1\leq i\neq j\leq N}$. Both nodes and edges are categorical variables, where $x^{(n)}\in\{1,\ldots,X\}$ and $e^{(i,j)}\in\{1,\ldots,E\}$.
@@
Graph generative models aim to learn the underlying probability distribution over graphs that produced a given dataset, enabling the generation of new samples that preserve the structural patterns and properties of the training graphs. Recently, Discrete Flow Matching (DFM)-based models gat2024discrete ; campbell2024generative have achieved state-of-the-art performance in undirected graph generation qin2025defog . DFM belongs to a broader class of discrete state-space methods that recover the original graph distribution $\bm{p}_{1}$ by progressively denoising samples from a pre-specified noise distribution $\bm{p}_{\text{noise}}$ vignac2023digress ; xu2024disco ; siraudin2024cometh . By operating directly in the discrete spaces, these methods naturally align with the discreteness of graphs and are well-suited to capturing complex dependencies through iterative refinement. DFM distinguishes itself from other methods in this class by decoupling sampling from training, allowing for post-training sampling optimization. This additional layer of expressivity is particularly valuable for digraphs, where edge directionality introduces combinatorial complexity and exponentially expands the search space. We now introduce the two main processes of DFM: the noising and denoising processes.
@@
We use a subscript $t$ to denote variables at time $t$, e.g., $x_{t}^{(n)}$ or $e_{t}^{(i,j)}$. The noising process then runs from $t=1$ to $t=0$, progressively corrupting the original input graphs through a linear interpolation between the data distribution and the limit distribution $\bm{p}_{0}=\bm{p}_{\text{noise}}$. This interpolation is applied independently to each variable gat2024discrete ; campbell2024generative , corresponding to each node and edge in the case of graphs. For example, for each node $x_{1}^{(n)}$, its noisy distribution at step $t$ is defined as:
@@
where $\delta(\cdot,\cdot)$ denotes the Kronecker delta and $p^{X}_{\text{noise}}$ is the limit distribution over node categories. A similar construction is used for the edges qin2025defog .
@@
The denoising process aims to reverse the noising trajectory, running from $t=0$ to $t=1$. It leverages a denoising neural network, parametrized by $\theta$, which is trained to predict the clean node $\bm{p}_{1\mid t}^{\theta,(n)}$ and edge $\bm{p}_{1\mid t}^{\theta,(i,j)}$ distributions given a noisy graph, aggregated as:
@@
where the expectation is taken over time $t$, sampled from a predefined distribution over $[0,1]$ (e.g., uniform); $G_{1}\sim\bm{p}_{1}(G_{1})$ is a clean graph from the dataset; and $G_{t}\sim\bm{p}_{t}(G_{t}|G_{1})$ is its noised version at time $t$ qin2025defog . The hyperparameter $\lambda\in\mathbb{R}^{+}$ controls the relative weighting between node and edge reconstruction losses. The denoising process is then formulated as a Continuous-Time Markov Chain (CTMC), which starts from sampling a noisy graph $G_{0}\sim\bm{p}_{0}$ and evolves according to:
where $\bm{R}_{t}$ is the CTMC rate matrix campbell2024generative . In practice, this update rule is approximated in two ways: first, by discretizing time with a finite step size $\Delta t$, in an Euler method step; and second, by approximating the ground-truth rate matrix with an estimate $\bm{R}^{\theta}_{t}$ computed from the denoising network predictions $\bm{p}^{\theta}_{1\mid t}(\cdot\mid G_{t})$. Further details are provided in Appendix I.
@@ 3.1 Directed Graph Generation via Discrete Flow Matching
Building on the formulation of Sec. 2, we extend DFM to the directed graph setting. Inspired by standard practice in iterative refinement for undirected graphs vignac2023digress ; xu2024disco ; siraudin2024cometh ; qin2025defog , we assume that every ordered pair of distinct nodes corresponds to a directed edge belonging to one of several possible classes, including a class representing the absence of an edge. We provide the complete training and sampling algorithms in Appendix D.
To mitigate the combinatorial explosion induced by the asymmetric adjacency matrices in digraphs (see Figure 1(a)), we leverage a key advantage of DFM: its decoupling of training and sampling. This flexibility allows for post-training sampling optimization, including time-adaptive schedules and custom CTMC rate matrices qin2025defog , which we extend to the directed setting. Although these optimizations yield consistent improvements (see Sec. H.5), they alone are insufficient to fully capture the distinctive distributional and structural properties of directed graphs without incorporating additional inductive biases, as demonstrated in Section 5. To overcome this, we introduce architectural components explicitly designed to model these unique characteristics and enable effective directionality-aware graph generation.
@@ 3.2 Asymetric positional encoding
The generative framework introduced above employs a denoising GNN. However, GNNs inherently struggle to capture global or higher-order patterns due to the locality of message passing xu2019GIN ; morris2019goneural . To maximize the capacity of our method, we augment our GNN with positional encodings (PEs) that inject structural information beyond local neighborhoods beaini2021DirGraphNet ; bouritsas2022improving . Common choices for undirected graphs include Laplacian eigenvectors or shortest-path distances, which have proven effective in enriching node and edge representations for undirected graph generation vignac2023digress ; qin2025defog . However, these encodings do not respect edge directionality, and thus fail to distinguish asymmetric structural roles. To overcome this limitation, we adopt recent direction-aware positional encodings geisler2023transformers ; huang2025good that account for both incoming and outgoing connectivity. We append these encodings to node and edge features, allowing our model to better capture the structural dependencies unique to directed graphs.
Magnetic Laplacian
In the undirected case, eigenvectors of the combinatorial Laplacian (Lap) rely on the spectral decomposition $\bm{L}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{-1}$ where $\bm{L}$ is the combinatorial Laplacian, $\bm{\Lambda}$ the diagonal matrix of eigenvalues, and $\bm{\Gamma}$ the matrix of eigenvectors. In the directed case, the combinatorial Laplacian can be defined as $\bm{L}_{\text{dir}}=\bm{D}_{s}-\bm{A}_{s}$, where $\bm{A}_{s}$ is the symmetrized adjacency matrix and $\bm{D}_{s}$ the symmetrized degree matrix. This symmetrization discards directionality, as it enforces $\bm{A}_{s}=\bm{A}_{s}^{\top}$, but it is necessary to guarantee that $\bm{L}_{\text{dir}}$ is symmetric and positive semi-definite.
To retain directional information while preserving desirable spectral properties, the authors in geisler2023transformers propose the Magnetic Laplacian (MagLap), which introduces a complex-valued phase encoding into the adjacency matrix. It is given by:
$\bm{L}_{\text{dir}}^{(q)}=\bm{D}_{s}-(\bm{A}_{s}\odot\exp(i\Theta^{(q)}))$ (5)
where $\odot$ denotes the element-wise (Hadamard) product, $i$ is the imaginary unit, and the phase matrix $\Theta^{(q)}$ is defined as $\Theta^{(q)}_{u,v}=2\pi q(\bm{A}_{u,v}-\bm{A}_{v,u})$. The parameter $q\geq 0$ controls the strength of the phase shift. When $q=0$, the Magnetic Laplacian reduces to the classical combinatorial Laplacian. The resulting complex-valued eigenvectors can effectively encode directionality patterns in the graph, and we leverage them by concatenating the real and complex parts of the eigenvectors to the node features and the eigenvalues to the global graph features.
Multi-$q$ Magnetic Laplacian
Recent work by huang2025good introduces a multi-$q$ extension of the Magnetic Laplacian, which leverages $Q$ distinct complex-valued potentials $q_{1},\ldots,q_{Q}$ to recover a more informative bidirectional walk profile, effectively capturing asymmetries in node connectivity. By extending the notion of the walk profile to the directed setting, this method enhances the expressiveness of the Magnetic Laplacian. In our model, we leverage this information by concatenating the first $k$ eigenvalues (real and imaginary parts) of each of the $Q$ magnetic Laplacians to the global features, and the corresponding first $k$ eigenvectors to the node features, providing direction-aware structural signals to both levels of representation.
Directed Relative Random Walk Probabilities (RRWP)
For undirected graphs, RRWP encodes the likelihood of arriving from one node to another through $k$-step random walks, expressed via a transition matrix $\bm{T}^{k}=(\bm{A}\bm{D}^{-1})^{k}$. For digraphs, authors in geisler2023transformers adapt this idea by defining $\bm{T}=\bm{A}\bm{D}^{-1}_{out}$, using the degree matrix of the outgoing edges $\bm{D}^{-1}_{out}$. To account for reverse flows, they also suggest that it is informative to additionally model reverse random walks, representing the likelihood of reaching a given edge starting from another through $k$-step random walks with $\bm{R}=\bm{A}^{\top}\bm{D}_{\text{in}}^{-1}$, where $\bm{D}_{\text{in}}$ is the in-degree matrix.
Following geisler2023transformers , we ensure that the probabilities encoded in the matrices sum up to 1 by adding self-loops to sink nodes during pre-processing. Finally, the full positional encoding concatenates both forward and reverse walk features for the $K$-step transition probabilities, producing the PE:
$\operatorname{RRWP}(G)=[\bm{I},\bm{T},\bm{T}^{2},\ldots,\bm{T}^{K-1},\bm{I},\bm{R},\bm{R}^{2},\ldots,\bm{R}^{K-1}]$ (6)
that we concatenate to the edge features, where the diagonal elements of each matrix are used to enrich the node features. To mitigate the tendency of $k$-step random walks to concentrate on sink nodes for large $k$, we incorporate Personalized PageRank (PPR) features. The PPR matrix is defined in closed form as $\text{PPR}=p_{r}(\bm{I}-(1-p_{r})\bm{T})^{-1}$, where $p_{r}$ denotes the restart probability.
@@
We hypothesize that incorporating directionality-aware positional encodings improves the generative performance of digraphs. To test this, we design an ablation study that evaluates which of the described encodings provide the most informative structural signals for directed graph generation tasks. See Appendix B for a more detailed description of the positional encodings.

### v2

@@ Abstract
Directed graphs naturally model systems with asymmetric, ordered relationships, essential to applications in biology, transportation, social networks, or visual understanding. Generating such graphs enables simulation, data augmentation and novel instance discovery; however, this task remains underexplored. We identify two key reasons: first, modeling edge directionality introduces a substantially larger dependency space, making the underlying distribution harder to learn; second, the absence of standardized benchmarks hinders rigorous evaluation. Addressing the former limitation requires more expressive models that are sensitive to directional topologies. Thus, we propose Directo, the first generative model for directed graphs built upon the discrete flow matching framework. Our approach combines: (i) a dual-attention mechanism distinctly capturing incoming and outgoing dependencies, (ii) a robust, discrete generative framework, and (iii) principled positional encodings tailored to asymmetric pairwise relations. To address the second limitation and support evaluation, we introduce a novel and extensive benchmark suite covering synthetic and real-world datasets. Experiments show that our method outperforms existing directed graph generation approaches across diverse settings and competes with specialized models for particular classes, such as directed acyclic graphs. These results highlight the effectiveness and generality of our approach, establishing a solid foundation for future research in directed graph generation.
@@ 1 Introduction
Directed graphs (digraphs) model systems with asymmetric relationships, capturing essential structures such as flows, dependencies, and hierarchies that arise in many real-world applications. This makes digraphs particularly well-suited for problems in diverse domains including biology (Li et al., 2006; Takane et al., 2023; Wei et al., 2024), transportation (Concas et al., 2022), social dynamics (Schweimer et al., 2022), and, more recently, image and video understanding (Chang et al., 2023; Rodin et al., 2023), where structured and directional representations are critical for interpretation and reasoning. Consequently, generating digraphs is central to tasks such as simulation, data augmentation and novel instance discovery in domains with directional structure.
@@
Figure 1: (a) The learnable space increases drastically with the number of nodes for digraphs compared to undirected graphs (Harary & Palmer, 1973), highlighting challenges in extending graph generative models to directed structures. (b) Overview of our generative model for directed graphs. During training, the dual attention transformer (denoising network), parameterized by $\theta$, is enhanced with asymmetric positional encoding, learning to reverse predictions using cross-entropy loss. During inference, we compute the rate matrix $R_{t}^{\theta}(G_{t},G_{t+\Delta t})$, which governs the evolution of the generative process over finite intervals $\Delta t$, based on the model prediction $p^{\theta}(\cdot\mid G_{t})$.
@@
Graph generative models have shown strong potential in diverse applications, including drug discovery (Mercado et al., 2020; Vignac et al., 2023b), finance (Li et al., 2023), social network modeling (Tsai et al., 2023), and medicine (Nikolentzos et al., 2023). However, most generative models focus on undirected graphs (Liao et al., 2019; Martinkus et al., 2022; Vignac et al., 2023a; Siraudin et al., 2024; Xu et al., 2024; Eijkelboom et al., 2024; Qin et al., 2025a), with only limited efforts on directed settings where recent works propose auto-regressive models for DAGs (Zhang et al., 2019; Li et al., 2025) and, more recently, for general digraphs (Law et al., 2025). We identify two main factors limiting progress in this direction. First, at the modeling level, the directed setting is inherently more challenging than its undirected counterpart, as edge directionality greatly enlarges the learnable space (see Figure 1(a): for graphs with $4$ nodes, there are $218$ possible digraphs, compared to only $11$ undirected ones). We will see that simply extending undirected architectures without explicit directional components fails on this larger space of digraphs. This is further amplified in settings where domain-specific structural constraints must be respected, such as directed acyclic graphs (DAGs). Second, beyond these technical challenges, we identify the lack of standardized benchmarks for directed graph generation as a barrier to rigorous evaluation and fair comparison. Current DAG methods (Li et al., 2025) typically rely only on proxy metrics derived from downstream tasks, with no established benchmarks to evaluate digraph generative quality systematically.
@@
To address the first challenge, we introduce Directo  [^1] Code available at: https://anonymous.4open.science/r/DirectoAnonymous, the first iterative refinement-based generative model for directed graphs. Our approach brings together three core components of expressive graph generative modeling: (i) an expressive neural architecture, (ii) a robust generative framework, and (iii) informative input features. For the architecture, we employ a dual attention mechanism that performs cross-attention between edge features and their reversed counterparts, capturing both source-to-target and target-to-source information flow. All this is integrated within a robust discrete-state flow matching framework that enables efficient generation with state-of-the-art performance and is also readily extendable to conditional generation tasks through classifier-free guidance (Ho & Salimans, 2022). Finally, the network’s input features include positional encodings designed for asymmetric adjacency matrices, which we show outperform directionality-agnostic alternatives. Figure 1(b) gives an overview of our method and its components.
@@
To tackle the second challenge, we propose a new benchmarking framework for rigorous evaluation of directed graph generation, spanning both synthetic and real-world datasets. The synthetic benchmarks cover different distributions, including acyclic and community-based graphs, while real-world benchmarks include a DAG dataset for Neural Architecture Search (NAS) (Phothilimthana et al., 2023) and a scene understanding dataset (Krishna et al., 2017). The evaluation metrics are systematic and tailored to directed graphs, including distributional distances for fidelity, and dataset-specific metrics reflecting the graph types used in different tasks. We evaluate Directo on the proposed benchmarks, where it consistently outperforms prior approaches across all settings, underscoring its effectiveness on structurally diverse datasets. Extensive ablations further reveal that the dual attention mechanism is critical for modeling directional dependencies, while positional encodings play a role in enhancing overall generation quality.
@@
    We propose Directo, the first flow-based digraph generative model, combining a direction-aware dual attention block with positional encodings that capture asymmetric structure.
@@
    We address the lack of benchmarks by releasing an evaluation suite covering synthetic graphs with relevant structural properties and real-world datasets with meaningful applications.
@@
    We conduct an extensive empirical analysis demonstrating that Directo achieves state-of-the-art performance on structurally diverse synthetic and real-world graphs, while showing the need for a dedicated directed architecture.
Overall, by addressing the two main bottlenecks in directed graph generation, we propose an effective, new generative framework and provide standardized benchmarks, aiming to lay the groundwork for future research and to support real-world applications.
@@ 2 Background: Discrete Flow Matching for Graph Generation
We denote by $G=\left(x^{(1:n:N)},\,e^{(1:i\neq j:N)}\right)$ directed graphs with $N$ nodes. The set of nodes is denoted by $\{x^{(n)}\}_{n=1}^{N}$, and the set of directed edges by $\{e^{(i,j)}\}_{1\leq i\neq j\leq N}$. Both nodes and edges are categorical variables, where $x^{(n)}\in\{1,\ldots,X\}$ and $e^{(i,j)}\in\{1,\ldots,E\}$. Inspired by standard practice in iterative refinement for undirected graphs (Vignac et al., 2023a; Xu et al., 2024; Siraudin et al., 2024; Qin et al., 2025a), we assume that every ordered pair of distinct nodes corresponds to a directed edge in one of several possible classes, including a class representing an absent edge. Finally, we use a subscript $t$ to denote variables at time $t$, e.g., $x_{t}^{(n)}$ or $e_{t}^{(i,j)}$.
@@
Graph generative models aim to learn the probability distribution of a set of observed graphs, enabling the generation of new samples that preserve structural and statistical properties. Discrete Flow Matching (DFM)-based models (Gat et al., 2024; Campbell et al., 2024) have recently achieved state-of-the-art performance in undirected graph generation (Qin et al., 2025a). DFM belongs to a broader class of discrete state-space methods that recover the original graph distribution $\bm{p}_{1}$ by progressively denoising samples from a pre-specified noise distribution $\bm{p}_{\text{noise}}$ (Vignac et al., 2023a; Xu et al., 2024; Siraudin et al., 2024). It distinguishes itself from other methods in this class by decoupling sampling from training, allowing for post-training sampling optimization. Additionally, by operating directly in the discrete spaces, these methods naturally align with the discreteness of graphs and are well-suited to capture complex dependencies through iterative refinement. We hypothesize that this added expressivity is especially valuable for modeling the added complexity of digraphs. We now introduce the two main processes of DFM: noising and denoising.
@@
The noising process runs from $t=1$ to $t=0$, progressively corrupting the original input graphs through a linear interpolation between the data distribution and the limit distribution $\bm{p}_{0}=\bm{p}_{\text{noise}}$ (as detailed in Figure 1(b)). This interpolation is applied independently to each variable (Gat et al., 2024; Campbell et al., 2024), corresponding to each node and edge in the case of graphs. For example, for each node $x_{1}^{(n)}$, its noisy distribution at step $t$ is defined as:
@@
where $\delta(\cdot,\cdot)$ denotes the Kronecker delta and $p^{X}_{\text{noise}}$ is the limit distribution over node categories. A similar construction is used for the edges (Qin et al., 2025a).
@@
The denoising process aims to reverse the noising trajectory, running from $t=0$ to $t=1$. It is formulated as a Continuous-Time Markov Chain (CTMC), which starts sampling a noisy graph $G_{0}\sim\bm{p}_{0}$ and evolves according to:
where $\bm{R}_{t}$ is the CTMC rate matrix (Campbell et al., 2024). In practice, this update rule is approximated in two ways: first, by discretizing time with a finite step size $\Delta t$, in an Euler method step; and second, by approximating the ground-truth rate matrix with an estimate $\bm{R}^{\theta}_{t}$ computed from the predictions of a denoising neural network $\bm{p}^{\theta}_{1\mid t}(\cdot\mid G_{t})$ parametrized by $\theta$. This network is trained to predict the clean node $\bm{p}_{1\mid t}^{\theta,(n)}$ and edge $\bm{p}_{1\mid t}^{\theta,(i,j)}$ distributions given a noisy graph, aggregated as:
@@
where the expectation is taken over time $t$, sampled from a predefined distribution over $[0,1]$ (e.g., uniform); $G_{1}\sim\bm{p}_{1}(G_{1})$ is a clean graph from the dataset; and $G_{t}\sim\bm{p}_{t}(G_{t}|G_{1})$ is its noised version at time $t$ (Qin et al., 2025a). The hyperparameter $\lambda\in\mathbb{R}^{+}$ controls the relative weighting between node and edge reconstruction losses. Further details are provided in Appendix I.
@@ 3.1 Directed Graph Generation via Discrete Flow Matching
Building on the formulation of Section 2, we extend DFM to the directed graph setting. To address the added complexity of asymmetric adjacency matrices in digraphs (Figure 1(a)), we exploit a key strength of DFM: the decoupling of training and sampling. This enables post-training optimizations such as time-adaptive schedules and custom CTMC rate matrices (Qin et al., 2025a), which we can extend to directed graphs (details in Appendix C). While these strategies improve performance (see Section H.8), they remain insufficient to capture the structural properties of digraphs (see Section 5). We therefore introduce architectural components explicitly tailored for directionality-aware graph generation, providing the complete training and sampling algorithms in Appendix D.
@@ 3.2 Asymetric positional encoding
Our generative framework employs a denoising GNN, which inherently struggles to capture global or higher-order patterns due to the locality of message passing (Xu et al., 2019; Morris et al., 2019). To maximize the capacity of our method, we augment our GNN with positional encodings (PEs) that inject structural information beyond local neighborhoods (Beaini et al., 2021; Bouritsas et al., 2020). Common choices for undirected graphs include Laplacian eigenvectors or shortest-path distances, effective in enriching node and edge representations (Vignac et al., 2023a; Qin et al., 2025a). However, these encodings overlook edge directionality, failing to distinguish asymmetric structural roles. To overcome this, we adopt direction-aware positional encodings (Geisler et al., 2023; Huang et al., 2025) that account for both outgoing and incoming connectivity. We append them to node and edge features, allowing our model to better capture the structural dependencies unique to digraphs.
Specifically, we experiment with three families of encodings: (i) the Magnetic Laplacian (MagLap) (Geisler et al., 2023), which introduces complex-valued phase shifts into the standard Laplacian to retain edge orientation; (ii) its Multi-$q$ MagLap extension (Huang et al., 2025), which stacks $Q$ Multiple Magnetic Laplacians with different complex potentials to provide richer representations; and (iii) Directed Relative Random Walk Probabilities (RRWP) (Geisler et al., 2023), which combine forward and reverse transition probabilities to capture outgoing and incoming asymmetric flows. These encodings are designed to provide direction-sensitive structural signals that better reflect the properties of directed graphs, and are described in detail in Appendix B.
@@
We hypothesize that incorporating directionality-aware positional encodings improves digraph generative performance. To test this, we design an ablation study evaluating which of the described encodings provide the most informative structural signals, while also considering their computational cost, to identify the best tradeoff between performance and efficiency.